find the domain of the function. g(x)= 6 / 3-7x
Answer by dabanfield(803)

find the domain of the function. g(x)= 6 / 3-7x
The function g is defined for all values of x except where the denominator is not defined (that is, not equal to 0).
The denominator 3-7x = 0 when x = 3/7 so the domain is all real nunmers except 3/7.